OpenStudy (anonymous):

Suppose that f(x)=sqrt (x+8), g(x)=x^2+7,and h(x)=x+9. what is f of g of h

OpenStudy (anonymous):

\[g(h(x)) = (x+9)^2 + 7\] \[f(g(h(x))) = \sqrt{((x+9)^2 + 7)+8}\]